You are planning to prepare “salami-and-cheese sandwiches” for your friends. You need to slice the salami you bought for this purpose, and you will buy cheese in grams and bread in packages, each containing 24 loaves which are approximately 10 cm x 10 cm squares. You have adjusted the food slicer to cut 2.0 mm thick slices from the 0.18 m long salami you bought that is 100 mm in diameter.

a)  How many bread packages do you need to buy if each “salami-and-cheese sandwich” contains three slices of salami? (Hint: Your limiting reagent is salami)

b) You decided to add one 2.0 mm thick cheese slice to each sandwich from a cheese block which is a prism with a 3.5 in x 3.5 in square base. How much cheese (in grams) do you need to buy if its density is 1.093 g/cm3?


1
Expert's answer
2022-03-01T17:46:01-0500

a. (10×10)cm2 = 100cm2

100×24=2400

b. (3.5×3.5) = 12.25